OpenStudy (anonymous):

PLEAE HELP I GIVE ANYTHING Using the completing-the-square method, find the vertex of the function f(x) = 5x2 + 10x + 8 and indicate whether it is a minimum or a maximum and at what point. Maximum at (1, 8) Minimum at (1, 8) Maximum at (–1, 3) Minimum at (–1, 3)

OpenStudy (anonymous):

@assatmath

OpenStudy (anonymous):

help please

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=5x^2+10x+8 }\) factor the 1st 2 terms out of 5, \(\large\color{black}{ \displaystyle f(x)=5(x^2+2x)+8 }\) then, tell me what number would you want added into the parenthesis ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

b/2squared right?

OpenStudy (solomonzelman):

yes, 1 is the number you need added. And now I will use a technique to which I refer to as a magic zero.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=5(x^2+2x)+8 }\) \(\large\color{black}{ \displaystyle f(x)=5(x^2+2x\color{blue}{+1}\color{red}{-1})+8 }\)

OpenStudy (solomonzelman):

then, take the negative one (which I labelled in red) out of the parenthesis, by expanding.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=5(x^2+2x\color{blue}{+1}\color{red}{-1})+8 }\) \(\large\color{black}{ \displaystyle f(x)=5(x^2+2x\color{blue}{+1})+5(\color{red}{-1})+8 }\) \(\large\color{black}{ \displaystyle f(x)=5(x^2+2x\color{blue}{+1})-5+8 }\) \(\large\color{black}{ \displaystyle f(x)=5(x^2+2x\color{blue}{+1})+3 }\)

OpenStudy (solomonzelman):

Now, you need to do the very last step

OpenStudy (anonymous):

5(x+1)^2+3?

OpenStudy (anonymous):

answer choice D?

OpenStudy (solomonzelman):

yes